forces 5N, 12N and 13N acting on a body are in equilibrum. What is the angle between the forces 5N and 12N
1
Expert's answer
2020-05-14T09:18:19-0400

Solution.

"F_1=5N;"

"F_2=12N;"

"F_3=13N;"

"\overrightarrow{F_1}+\overrightarrow{F_2}+\overrightarrow{F_3}=0;"

"\overrightarrow{F_1}+\overrightarrow{F_2}=\overrightarrow{F_3};"

f the forces are perpendicular, then we find the equivalent of the two forces by Pythagoras' theorem, otherwise we will use the cosine theorem.

"\overrightarrow{F_{12}}=\overrightarrow{F_1}+\overrightarrow{F_2};"

First, check whether the forces are perpendicular:

"F_{12}=\sqrt{F_1^2+F_2^2};"

"F_{12}=\sqrt{(5N)^2+(12N)^2}=13N;"

"F_{12}=F_3;"

Therefore, the forces 5N and 12N form an angle of 90o;

Answer: the forces 5N and 12N form an angle of 90o.